Doubling Time: If interest is compounded continuously and the interest rate is tripled, what effect will this have on the time required for an investment to double?
***I know that we are suppose to use the equation:
A = P(e^(rt))

Answer by stanbon(75887)

Doubling Time: If interest is compounded continuously and the interest rate is tripled, what effect will this have on the time required for an investment to double?
-------------
Before tripling:
2Ao = Ao*e^(rt)
2 = e^rt
rt = ln(2)
t = (1/r)ln2
-------------------
After tripling:
2Ao = Ao*e^(3rt)
2 = e^(3rt)
3rt = ln2
t = (1/3r)ln2
================
Cheers,
Stan H.